iamchris08:

Rewrite the radical expression as an expression with rational exponents. fourth root of x to the power of 5

dude:

Lets first write out what they gave us: \(\large \sqrt[4]{x^5}\) The rule you should note about roots and fraction exponents is: \(\large \sqrt[n]{x^m}=x^{\frac{m}{n}}\) [The exponent goes on the numerator and the root is on the denominator]